ganeshie8 (ganeshie8):

eliminate \(v\) \(p = \dfrac{mv}{\sqrt{1-(v/c)^2}}\) \(K=mc^2\left(\dfrac{1}{\sqrt{1-(v/c)^2}}-1\right)\)

ganeshie8 (ganeshie8):

this is similar to eliminating v from the momentum, KE equations and getting the relation p^2/2m = ke

ganeshie8 (ganeshie8):

eliminate \(v\) \(p = \dfrac{mv}{\sqrt{1-(v/c)^2}}\) \(K=mc^2\left(\dfrac{1}{\sqrt{1-(v/c)^2}}-1\right)\)

Parth (parthkohli):

\[\left(\frac{K}{mc^2} +1\right)^{-2} = 1 - (v/c)^2\]So on to solve for \(v\) and substituting in the first equation.

Parth (parthkohli):

I don't see a good form for this though :|

ganeshie8 (ganeshie8):

Nice! \[1-\left(\frac{K}{mc^2} +1\right)^{-2} = (v/c)^2\]

ganeshie8 (ganeshie8):

im thinking of squaring the first equation and substituting above thingy

Parth (parthkohli):

Oh, yeah, that'd work I guess.

ganeshie8 (ganeshie8):

\[p^2 = \dfrac{(mv)^2}{1-(v/c)^2} \implies (p/c)^2 = \dfrac{m^2(v/c)^2}{1-(v/c)^2} \]

ganeshie8 (ganeshie8):

that still looks nasty... somehow i need to arrive at below relation : \[(pc)^2=K^2+2Kmc^2\]

Parth (parthkohli):

can we say something like\[\left(1 + \frac{K}{mc^2} \right)^{-2} \approx 1 - 2\frac{K}{mc^2}\]this only works if \(K \ll mc^2\) and I'm not really sure if that's true in relativistic mechanics haha.

ganeshie8 (ganeshie8):

approximating using taylor series ?

Parth (parthkohli):

yes, but again that only works if \(K \ll mc^2\).

ganeshie8 (ganeshie8):

kinetic energy can be a million times greater than the rest energy mc^2, so that approximation is not so useful here..

ganeshie8 (ganeshie8):

for example rest energy (mc^2) of an electron is just 0.511 MeV, but we can accelerate the electrons to huge kinetic energies as 50 GeV

Parth (parthkohli):

that's weird. as long as you're not doing anything wrong, you should be able to get that expression. and yeah, I just looked at the original expression for kinetic energy and it makes sense how for low values of velocity it's way bigger than \(mc^2\)

ganeshie8 (ganeshie8):

yeah this looks messy... il just substitute K and p in the final equation and verify if it balances..

Parth (parthkohli):

I just looked at your expression and seeing some patterns\[(pc)^2 = K^2 + 2K mc^2+ \color{#C00}{(mc^2)^2 - (mc^2)^2} \]

Parth (parthkohli):

as you said the kinetic energy can be a million times greater than the rest energy. it's possible that the actual expression is\[(pc)^2 = K^2 + 2K mc^2 + (mc^2)^2 = (K + mc^2)^2\]and that they neglected the final term. wait, let me check if this is true. :P

Parth (parthkohli):

nah im dumb

ganeshie8 (ganeshie8):

that looks very interesting! we can't give away K or mc^2 though... their relative values depend on the mass of the particles...

ganeshie8 (ganeshie8):

for an electron, rest energy may be negligible compared to the kinetic energy but for an uranium atom, rest energy may not be negligibel..

Parth (parthkohli):

hey did you plug in the values of K and p in the equation you're given in the end? i think plugging in can have two benefits: it can tell us if an equation is exactly true, and if not, we can simplify it and end up with something in the form \(\alpha = \beta\) which would mean that we've approximated \(\alpha\) to be equal to \(\beta\) if you know what i mean. basically we're reverse-solving it.

ganeshie8 (ganeshie8):

i haven't plugged in p and K yet, but i plan to do it as soon as i find a paper and pen... here is a screenshot from textbook :

ganeshie8 (ganeshie8):

looks the textbook is not referring to any approximation...